# An Observability-Constrained Magnetic-Field-Aided Inertial Navigation System — Extended Version

### v1

@@ Abstract
A method to construct an observability-constrained magnetic-field-aided inertial navigation system is proposed. The proposed method builds upon the previously proposed observability-constrained extended Kalman filter and extends it to work with a magnetic-field-based odometry-aided inertial navigation system. The proposed method is evaluated using simulation and real-world data, showing that (i) the system observability properties are preserved, (ii) the estimation accuracy increases, and (iii) the perceived uncertainty calculated by the EKF is more consistent with the true uncertainty of the filter estimates.
@@ I Introduction
In an odometry-aided inertial navigation system (INS), such as a magnetic-field odometry-aided INS [1, 2, 3, 4, 5, 6], the uncertainty about the position and yaw of the navigation platform can, if the errors in the initial state are uncorrelated, never become smaller than the initial position and yaw uncertainty. This is because odometry and inertial measurements only provide relative motion information. Hence, for any algorithm used to estimate the navigation state $x_{k}$ in an odometry-aided INS it should hold that
@@
In [1], an extended Kalman filter (EKF) algorithm was used to realize a magnetic-field-aided INS (MAINS). In Fig. 1, the square roots of $P^{\phi}_{0}$ and $P^{\phi}_{k}$ as calculated by the EKF in the MAINS are shown. As seen from the figure the square root of $P^{\phi}_{k}$ falls below that of $P^{\phi}_{0}$. Hence the inequality in (1) does not hold and the covariance estimate of the EKF is inconsistent. The inconsistency is because the EKF linearizes the system model around the estimated state, which causes the yaw to be perceived as observable even though it is not. This type of inconsistency effect has been observed in multiple EKF-based implementations of odometry-aided INS, see e.g., [7, 8, 9], and is troublesome if the yaw information should be fused with information from other systems or be used in a control or decision system. Next, we will present a way to fix this inconsistency by modifying the EKF algorithm so that it preserves the observability properties of the underlying nonlinear system. The presented method is an extension of the method published in [8]. Our contributions are:
@@
Figure 1: Example of the perceived yaw uncertainty calculated by the EKF used to realize magnetic-field based odometry aided INS in [1]. Also shown is the perceived yaw uncertainty of the proposed observability-constrained EKF algorithm.
@@ III Application to the MAINS
The MAINS proposed in [1] utilizes an IMU and an array of magnetometers, see Fig. 2, to perform odometry. The method presented in Section II will next be used to modify the EKF algorithm used in the MAINS to address the inconsistency illustrated in Fig. 1.
@@ III-A State-space Model
Consider a simplified MAINS system where IMU biases are not included in the state vector. The IMU biases are excluded as they only complicate the forthcoming observability analysis and biases never improve the observability. Later, when evaluating the proposed method the biases are included.
@@
Here the superscript n denotes the navigation frame where the physical quantity is resolved. Further, $p_{k}^{\text{n}}\in\mathbb{R}^{3},v^{\text{n}}_{k}\in\mathbb{R}^{3}$ denote the position and velocity, respectively. Moreover, $q_{k}\in\mathbb{H}$ denotes the unit quaternion that encodes the orientation of the body frame w.r.t. the navigation frame, and $\theta_{k}\in\mathbb{R}^{\kappa}$ denotes the coefficients of the polynomial model used to describe the magnetic field; see [1] for details.
@@
Recalling the notation used in [1], the state dynamics is given by
@@
$f(x_{k},u_{k},w_{k})$ $=\begin{bmatrix}p_{k}^{\text{n}}+v_{k}^{\text{n}}t_{\text{s}}+(R_{k}s_{k}+\text{g})\frac{t_{\text{s}}^{2}}{2}\\ v_{k}^{\text{n}}+(R_{k}s_{k}+\text{g})t_{\text{s}}\\ q_{k}\otimes\text{exp}_{q}\left(\omega_{k}t_{\text{s}}\right)\\ A^{\dagger}B(\psi_{k})\theta_{k}+w_{k}^{\theta}\end{bmatrix}.$ (9b)
Here $u_{k}\triangleq[s_{k}^{\top}\;\omega_{k}^{\top}]^{\top}$ is the system input, where $s_{k}\in\mathbb{R}^{3}$ and $\omega_{k}\in\mathbb{R}^{3}$ denote the specific force and angular velocity in body frame, respectively. $w_{k}^{\theta}\in\mathbb{R}^{\kappa}$ denotes the process noise affecting the coefficients of the magnetic field model. Furthermore, $t_{\text{s}}$ denotes the sampling period, $R_{k}\in SO(3)$ denotes the rotation matrix corresponding to $q_{k}$, $\otimes$ denotes quaternion multiplication, and $\text{exp}_{q}(\cdot)$ maps an axis-angle rotation to a quaternion. Moreover, the vector $\text{g}\in\mathbb{R}^{3}$ denotes the local gravity. The matrices $A^{\dagger}$ and $B(\cdot)$ are used in [1] to propagate coefficients of the magnetic field model; see [15] and [1] for details. Lastly, $\psi_{k}\in\mathbb{R}^{6}$ denotes the pose change between time $k$ and $k+1$, i.e.,
@@ III-B Linearized Error State Model
Here the standard additive error definition is used, except for the orientation error. For example, $\delta v_{k}^{\text{n}}=v^{\text{n}}_{k}-\bar{v}_{k}^{\text{n}}$). The orientation error $\epsilon_{k}\in\mathbb{R}^{3}$ satisfies the equation $R_{k}\approx\bar{R}_{k}(I_{3}+[\epsilon_{k}]_{\times})$, where $[\cdot]_{\times}$ is an operator that maps a vector in $\mathbb{R}^{3}$ to a skew-symmetric matrix such that $\left[\xi\right]_{\times}b=\xi\times b$.
@@
Here ${\tilde{s}_{k}\in\mathbb{R}^{3}}$ and ${\tilde{\omega}_{k}\in\mathbb{R}^{3}}$ denote the measured acceleration and angular velocity, respectively. Further, $w_{k}\triangleq[(w_{k}^{s})^{\top}\;(w_{k}^{\omega})^{\top}\;(w_{k}^{\theta})^{\top}]^{\top}$ denote the process noise, where $w_{k}^{s}\in\mathbb{R}^{3}$ and $w_{k}^{\omega}\in\mathbb{R}^{3}$ denote the acceleration and angular velocity measurement noise, respectively. The explicit form of $\bar{G}_{k}$ is not given as it is relevant in the observability analysis, interested readers can find it in [1].
@@ III-D Suggested Modifications of the Jacobians
The constraint (20b) does not concern the modification of the Jacobian, it is the result of keeping the first row of $\tilde{F}_{k}$ the same as that of $\hat{F}_{k}$.
@@ IV Experimental Evaluation
is used. Here $\rho$ denotes the quantity for which RMSE is computed, $M$ is the number of simulations, and $i$ denotes the the simulation index. Furthermore, the perceived uncertainty of the estimate at each timestep is calculated as
@@ IV-A Simulation Setup and Result
In the simulation, a sensor board, similar to that shown in Fig. 2, moves in squares of size 2 $\times$ 2 meters; see Fig. 3(a). The trajectory’s duration was 8 seconds and the data was sampled at 100 Hz. The IMU measurements were generated as the true value corrupted by additive white noise and biases. The magnetometer measurements were generated as the magnetic field from a multi-dipole model corrupted by additive white noise. The magnitude of the simulated magnetic field is in Fig. 3(a). In total, 50 simulations with independent noise, bias, and initial state realizations, were used in the Monte Carlo simulation evaluation. The same filter parameters were used in both algorithms.
The results are shown in Fig. 3(b) and Fig. 3(c). The figures show that the proposed OC-MAINS algorithm generally has a smaller position and yaw RMSE. Further, the perceived uncertainty of the yaw estimate is more consistent with the true uncertainty (See also Fig. 1.). Concerning the position estimates, even though the perceived uncertainty of the OC-MAINS is somewhat more consistent with the true uncertainty of the position estimates, significant inconsistency still exists. However, it is worth noting that the perceived uncertainty of OC MIANS is always higher than the initial uncertainty, thus satisfying the condition in (1). The original MAINS algorithm does not meet this condition.
@@ IV-B Experimental Setup and Result
In the experiment, a person held the sensor board in Fig. 2 parallel to the ground, and walked in squares for a few laps. The true trajectory was measured using a camera-based motion-tracking system. The same filter parameters were used in both algorithms.
The results are shown in Fig. 3(e) and Fig. 3(f). From Fig. 3(f) it can be seen that the yaw error of OC-MAINS is significantly lower than that of the original MAINS. Also, it can be seen that the perceived and true yaw uncertainty agree when using the OC-MAINS algorithm. From Fig. 3(e) it can be seen that the position error of the original MAINS and OC-MAINS algorithms are about the same, but the OC-MAINS algorithm has slightly better performance in the $y$-axis direction. For the OC-MAINS algorithm, the perceived and true uncertainty agrees in the $y$- and $z$-axis directions, whereas in the $x$-axis there is a significant inconsistency. This could result from imperfect IMU calibration. Similar to in the simulations, the perceived uncertainty of OC MIANS is always higher than the initial uncertainty, thus satisfying the condition in (1).
@@ V Summary & Conclusions
A method to construct an observability-constrained magnetic-field-aided INS has been proposed. The proposed method modifies the state transition matrix in the EKF to ensure that the observability properties of the nonlinear system are preserved in the filtering process. The method extends the work in [8] by introducing a transformation of the basis that spans the unobservable subspace. This transformation reduces the changes that must be done to the state transition matrix. The proposed method was evaluated on simulation and real-world datasets, showing that (i) the observability properties are preserved, (ii) the estimation accuracy increases, and (iii) the perceived uncertainty calculated by the EKF is more consistent with the true uncertainty of the filter estimates.
@@ Acknowledgment
This work has been funded by Security Link and the Swedish Research Council (Vetenskapsrådet) project 2020-04253 ”Tensor-field based localization”.
@@ Appendix A The Unobservable Subspace Basis
To derive the unobservable subspace basis for the system defined by (13) and (14), $\bar{F}_{k}$ is rewritten to its equivalent form. Propagating the nominal state $\bar{x}_{k}$ with $\tilde{u}_{k}$ using (9b) leads to
@@
Now, it can be verified that the columns of $N(\bar{x}_{k})$ in (15) are the basis vectors of the unobservable subspace by multiplying every block row in $\bar{\mathcal{O}}_{k}$ with $N(\bar{x}_{k})$. Let $\bar{\mathcal{O}}_{k}^{(s,l)}\triangleq H_{\delta x}^{(s)}\bar{\Phi}(l,k)$, where $H_{\delta x}^{(s)}$ denotes the block row in $H_{\delta x}$ corresponding to the $s$-th magnetometer. For $l=k$, $\bar{\Phi}(k,k)=I$, then
$\bar{\mathcal{O}}_{k}^{(s,k)}=\begin{bmatrix}0_{3\times 3}&0_{3\times 3}&0_{3\times 3}&H^{\theta}(r_{s})\end{bmatrix},$ (31)
and $\bar{\mathcal{O}}_{k}^{(s,k)}N(\bar{x}_{k})=\left[0_{3\times 3}\quad 0_{3\times 1}\right]$, which can be verified by simple calculation.
@@
$+\left.\>[\frac{2(j-k)+1}{2}t_{\text{s}}\text{g}]_{\times})\text{g}\right]$
@@
Since $\mathcal{\bar{O}}_{k}^{(s,l)}N(\bar{x}_{k})=0$ $\forall s,l\geq k$ then $\mathcal{\bar{O}}_{k}N(\bar{x}_{k})=0$ and the column vectors of $N(\bar{x}_{k})$ are the basis vectors of the unobservable subspace.
@@ Appendix B Interpretation of the Unobservable Subspace
$\delta x=c\begin{bmatrix}-[p^{\text{n}}]_{\times}\text{g}\\ -[v^{\text{n}}]_{\times}\text{g}\\ R^{\top}\text{g}^{\top}\\ 0\end{bmatrix}=c\left(-\begin{bmatrix}I_{3}\\ 0_{3\times 3}\\ 0_{3\times 3}\\ 0_{\kappa\times 3}\end{bmatrix}[p^{\text{n}}]_{\times}\text{g}+\begin{bmatrix}0\\ -[v^{\text{n}}]_{\times}\text{g}\\ R^{\top}\text{g}^{\top}\\ 0\end{bmatrix}\right)$ (40)

### v2

# An Observability-Constrained Magnetic Field-Aided Inertial Navigation System — Extended Version

@@ Abstract
Maintaining consistent uncertainty estimates in localization systems is crucial as the perceived uncertainty commonly affects high-level system components, such as control or decision processes. A method for constructing an observability-constrained magnetic field-aided inertial navigation system is proposed to address the issue of erroneous yaw observability, which leads to inconsistent estimates of yaw uncertainty. The proposed method builds upon the previously proposed observability-constrained extended Kalman filter and extends it to work with a magnetic field-based odometry-aided inertial navigation system. The proposed method is evaluated using simulation and real-world data, showing that (i) the system observability properties are preserved, (ii) the estimation accuracy increases, and (iii) the perceived uncertainty calculated by the EKF is more consistent with the true uncertainty of the filter estimates.
@@ I Introduction
Maintaining consistent uncertainty estimates in localization systems is crucial as the perceived uncertainty commonly affects high-level system components, such as control or decision processes. In an odometry-aided inertial navigation system (INS), such as the magnetic field-odometry-aided INSs presented in [1, 3, 4, 6], the uncertainty about the position and yaw of the navigation platform can, if the errors in the initial state are uncorrelated, never become smaller than the initial position and yaw uncertainty. This is because odometry and inertial measurements only provide relative motion information. Hence, for any algorithm used to estimate the navigation state $x_{k}$ in an odometry-aided INS it should hold that
@@
In [1], an extended Kalman filter (EKF) was used in a magnetic field-aided INS (MAINS). In Fig. 1, the square roots of $P^{\phi}_{0}$ and $P^{\phi}_{k}$ as calculated by the EKF in the MAINS are shown. As seen from the figure the square root of $P^{\phi}_{k}$ falls below that of $P^{\phi}_{0}$. Hence the inequality in (1) does not hold and the uncertainty estimate of the EKF is inconsistent. The inconsistency is because the EKF linearizes the system model around the estimated state, which causes the yaw to be perceived as observable even though it is not. This type of inconsistency effect has been observed in multiple EKF-based implementations of odometry-aided INS, see e.g., [7, 8, 9], and is troublesome if the yaw information should be fused with information from other systems or be used in a control or decision process. Next, we will present a way to fix this inconsistency by modifying the EKF algorithm so that it preserves the observability properties of the underlying nonlinear system. The presented method is an extension of the method published in [8]. Our contributions are:
@@
Figure 1: Example of the perceived yaw uncertainty calculated by the EKF used to realize magnetic field-odometry-aided INS in [1] (black line). Also shown is the perceived yaw uncertainty of the proposed observability-constrained EKF algorithm (red line), as well as the initial uncertainty (blue line).
@@ III Application to the MAINS
The MAINS utilizes an IMU and an array of magnetometers, see Fig. 2, to perform odometry. The method presented in Section II will next be used to modify the EKF algorithm used in the MAINS to address the inconsistency illustrated in Fig. 1. Due to space constraints, the system model presentation next does not include derivations or an in-depth description of the underlying physics. The interested reader is encouraged to read [1] for details.
@@ III-A State-Space Model
Consider a simplified MAINS system in which IMU biases are not included in the state vector. The biases are excluded as they only complicate the forthcoming observability analysis, and biases never improve observability. Later, when evaluating the proposed method, the biases are included.
@@
Here, the superscript n denotes the navigation frame where the physical quantity is resolved. Further, $p_{k}^{\text{n}}\in\mathbb{R}^{3},v^{\text{n}}_{k}\in\mathbb{R}^{3}$ denote the position and velocity, respectively. Moreover, $q_{k}\in\mathbb{H}$ denotes the unit quaternion that encodes the orientation of the body frame w.r.t. the navigation frame, and $\theta_{k}\in\mathbb{R}^{\kappa}$ denotes the coefficients of the polynomial model used to describe the magnetic field; see [1] for details.
@@
The state dynamics is given by [1]
@@
$f(x_{k},u_{k},w_{k}^{\theta})$ $=\begin{bmatrix}p_{k}^{\text{n}}+v_{k}^{\text{n}}t_{\text{s}}+(R_{k}s_{k}+\text{g})\frac{t_{\text{s}}^{2}}{2}\\ v_{k}^{\text{n}}+(R_{k}s_{k}+\text{g})t_{\text{s}}\\ q_{k}\otimes\text{exp}_{q}\left(\omega_{k}t_{\text{s}}\right)\\ A^{\dagger}B(\psi_{k})\theta_{k}+w_{k}^{\theta}\end{bmatrix}.$ (9b)
Here $u_{k}\triangleq[s_{k}^{\top}\;\omega_{k}^{\top}]^{\top}$ is the system input, where $s_{k}\in\mathbb{R}^{3}$ and $\omega_{k}\in\mathbb{R}^{3}$ denote the specific force and angular velocity in body frame, respectively. $w_{k}^{\theta}\in\mathbb{R}^{\kappa}$ denotes the process noise affecting the coefficients of the magnetic field model. Furthermore, $t_{\text{s}}$ denotes the sampling period, $R_{k}\in SO(3)$ denotes the rotation matrix corresponding to $q_{k}$, $\otimes$ denotes quaternion multiplication, and $\text{exp}_{q}(\cdot)$ maps an axis-angle rotation to a quaternion. Moreover, the vector $\text{g}\in\mathbb{R}^{3}$ denotes the local gravity. The matrices $A^{\dagger}$ and $B(\cdot)$ are used in [1] to propagate coefficients of the magnetic field model; see [1] for details. Lastly, $\psi_{k}\in\mathbb{R}^{6}$ denotes the pose change between time $k$ and $k+1$, i.e.,
@@ III-B Linearized Error State Model
Here the standard additive error definition is used, except for the orientation error. For example, $\delta v_{k}^{\text{n}}=v^{\text{n}}_{k}-\bar{v}_{k}^{\text{n}}$. The orientation error $\epsilon_{k}\in\mathbb{R}^{3}$ satisfies the equation $R_{k}\approx\bar{R}_{k}(I_{3}+[\epsilon_{k}]_{\times})$, where $[\cdot]_{\times}$ is an operator that maps a vector in $\mathbb{R}^{3}$ to a skew-symmetric matrix such that $\left[\xi\right]_{\times}b=\xi\times b$.
@@
Here ${\tilde{s}_{k}\in\mathbb{R}^{3}}$ and ${\tilde{\omega}_{k}\in\mathbb{R}^{3}}$ denote the measured acceleration and angular velocity, respectively. Further, $w_{k}\triangleq[(w_{k}^{s})^{\top}\;(w_{k}^{\omega})^{\top}\;(w_{k}^{\theta})^{\top}]^{\top}$ denote the process noise, where $w_{k}^{s}\in\mathbb{R}^{3}$ and $w_{k}^{\omega}\in\mathbb{R}^{3}$ denote the acceleration and angular velocity measurement noise, respectively. The explicit form of $\bar{G}_{k}$ is not given as it is relevant in the observability analysis; interested readers can find it in [1].
@@ III-D Suggested Modifications of the Jacobians
The constraint (20b) does not concern the modification of the Jacobian. It is the result of keeping the first row of $\tilde{F}_{k}$ the same as that of $\hat{F}_{k}$.
@@ IV Experimental Evaluation
is used. Here, $\rho$ denotes the quantity for which RMSE is computed, $M$ is the number of simulations, and $i$ denotes the simulation index. Furthermore, the perceived uncertainty of the estimate at each timestep is calculated as
@@ IV-A Simulation Setup and Result
In the simulation, a sensor board, similar to that shown in Fig. 2, moves in squares of size 2 $\times$ 2 meters; see Fig. 3(a). The trajectory’s duration was 8 seconds, and the data was sampled at 100 Hz. The IMU measurements were generated as the true value corrupted by white noise and biases. The magnetometer measurements were generated as the magnetic field from a multi-dipole model corrupted by additive white noise. The magnitude of the simulated magnetic field is in Fig. 3(a). In total, 50 simulations with independent noise, bias, and initial state realizations were used in the Monte Carlo simulation evaluation. The same parameter settings were used in both algorithms. Due to space constraints, all the settings are not presented here, but they can be found in the published code.
The results are shown in Fig. 3(b) and Fig. 3(c). The figures show that the proposed OC-MAINS algorithm generally has a smaller position and yaw RMSE. Further, the perceived uncertainty of the yaw estimate is more consistent with the true uncertainty (See also Fig. 1.). Concerning the position estimates, even though the perceived uncertainty of the OC-MAINS is somewhat more consistent with the true uncertainty of the position estimates, significant inconsistency still exists. However, it is worth noting that the perceived uncertainty of OC-MAINS is always higher than the initial uncertainty, thus satisfying the condition in (1). The original MAINS algorithm does not meet this condition.
@@ IV-B Experimental Setup and Result
In the experiment, a person held the sensor board in Fig. 2 parallel to the ground and walked in squares for a few laps. The true trajectory was measured using a camera-based tracking system. The same parameters were used in both algorithms.
The results are shown in Fig. 3(e) and Fig. 3(f). From Fig. 3(f), it can be seen that the yaw error of OC-MAINS is significantly lower than that of the original MAINS. Also, it can be seen that the perceived and true yaw uncertainty agree when using the OC-MAINS algorithm. From Fig. 3(e), it can be seen that the position error of the original MAINS and OC-MAINS algorithms are about the same, but the OC-MAINS algorithm has slightly better performance in the $y$-axis direction. For the OC-MAINS algorithm, the perceived and true uncertainty agrees in the $y$- and $z$-axis directions, whereas in the $x$-axis, there is a significant inconsistency. This could result from imperfect IMU calibration. Similar to in the simulations, the perceived uncertainty of OC-MAINS is always higher than the initial uncertainty, thus satisfying the condition in (1).
@@ V Summary & Conclusions
A method to construct an observability-constrained magnetic field-aided INS has been proposed. The proposed method modifies the state transition matrix in the EKF to ensure that the observability properties of the nonlinear system are preserved in the filtering process. The method extends the work in [8] by introducing a transformation of the basis that spans the unobservable subspace. This transformation reduces the changes that must be done to the state transition matrix. The proposed method was evaluated on simulation and real-world datasets, showing that (i) the observability properties are preserved, (ii) the estimation accuracy increases, and (iii) the perceived uncertainty calculated by the EKF is more consistent with the true uncertainty of the estimates.
@@ Acknowledgment
This work has been funded by Security Link and the Swedish Research Council project 2020-04253.
@@ Appendix A The Unobservable Subspace Basis
To derive the unobservable subspace basis for the system defined by (13) and (14), $\bar{F}_{k}$ is rewritten. Propagating the nominal state $\bar{x}_{k}$ with $\tilde{u}_{k}$ using (9b) leads to
@@
Now, it can be verified that the columns of $N(\bar{x}_{k})$ in (15) are unobservable error states by multiplying every block row in $\bar{\mathcal{O}}_{k}$ with $N(\bar{x}_{k})$. Let $\bar{\mathcal{O}}_{k}^{(s,l)}\triangleq H_{\delta x}^{(s)}\bar{\Phi}(l,k)$, where $H_{\delta x}^{(s)}$ denotes the block row in $H_{\delta x}$ corresponding to the $s$-th magnetometer. For $l=k$,
$\bar{\mathcal{O}}_{k}^{(k,k)}=\begin{bmatrix}0_{3\times 3}&0_{3\times 3}&0_{3\times 3}&H^{\theta}(r_{s})\end{bmatrix},$ (31)
and $\bar{\mathcal{O}}_{k}^{(s,k)}N(\bar{x}_{k})=\left[0_{3\times 3}\quad 0_{3\times 1}\right]$, which can be verified by simple calculation using that $\bar{\Phi}(k,k)=I$.
@@
$+\left.\>[\frac{2(j-k)+1}{2}t_{\text{s}}\text{g}]_{\times}\text{g})\right]$
@@
Since $\mathcal{\bar{O}}_{k}^{(s,l)}N(\bar{x}_{k})=0$ $\forall s,l\geq k$ then $\mathcal{\bar{O}}_{k}N(\bar{x}_{k})=0$ and the column vectors of $N(\bar{x}_{k})$ belong to the null space of $\bar{\mathcal{O}}_{k}$.
@@ Appendix B Interpretation of the Unobservable Subspace
$\delta x=c\begin{bmatrix}-[p^{\text{n}}]_{\times}\text{g}\\ -[v^{\text{n}}]_{\times}\text{g}\\ R^{\top}\text{g}\\ 0\end{bmatrix}=c\left(-\begin{bmatrix}I_{3}\\ 0_{3\times 3}\\ 0_{3\times 3}\\ 0_{\kappa\times 3}\end{bmatrix}[p^{\text{n}}]_{\times}\text{g}+\begin{bmatrix}0\\ -[v^{\text{n}}]_{\times}\text{g}\\ R^{\top}\text{g}\\ 0\end{bmatrix}\right)$ (40)